OpenStudy (anonymous):

Can you simplify square root of 148?

OpenStudy (anonymous):

148 = 4 * 37 SQRT(148) = SQRT(4) * SQRT(37) = 2 * SQRT(37)

OpenStudy (anonymous):

Oh, okay, Thank you. (:

OpenStudy (anonymous):



OpenStudy (anonymous):

what about square root of 32?

OpenStudy (anonymous):